# Robustness Analysis of USmorph: I. Generalization Efficiency of Unsupervised Strategies and Supervised Learning in Galaxy Morphological Classification

### v1

@@ 1 Introduction
2 Observation and Sample Seletction
@@
[…]
@@
3 Analysis for The Unsupervised Method for Morphological Classification
@@
3.2 Convolutional Autoencoder Framework for Noise Reduction
@@
3.3 The APCT Preprocessing Strategies
@@
Figure 5 illustrates the distribution of the mean pixel-value differences before and after a 90 rotation of galaxy images, with and without APCT. The results show that, with APCT, the majority of images exhibit minimal pixel-value differences, whereas without APCT, the differences are significantly larger due to the lack of perfect central symmetry in galaxies. This result demonstrates that the APCT effectively eliminates the effect of rotation on pixel distribution while preserving the core features of the image by converting Cartesian coordinates to polar coordinates. By transforming the problem of rotation invariance into one of translation invariance, the APCT can significantly enhance the rotational classification robustness of CNNs, which will be further validated in  5.1 via more tests.
3.4 Voting strategy and Post-hoc label alignment
@@
4 Analysis for The SML Models and the Experiment Settings
@@ 4.2 Experiment Setting
We further verify that key feature distributions were balanced between the two sets to ensure the model’s generalization capability and the reliability of performance evaluation, as well as to avoid potential biases caused by discrepancies in data distribution. As illustrated in Figure 7, the distributions of galaxy morphological types are consistent between the training and validation sets. Similarly, Figure 8 shows that their distributions in the parameter space of physical properties—such as stellar mass ($M_{\ast}$), redshift ($z$), and other morphological parameters are highly similar (see Section 5.3). The Kolmogorov–Smirnov (K–S; Kolmogorov+1933) test results (Table 1) show that all the examined parameters yield $p$-values greater than 0.05, indicating that the training and validation sets follow statistically consistent distributions.
@@
[…]
@@ 5.1 The effectiveness of CAE and APCT
5.2 Overall morphological classification results and t-SNE Visualization
@@
5.3 Comparisons of Galaxy Properties

### v2

@@ I Introduction
II Observation and Sample Seletction
@@
In addition to accurate photometric measurements, the COSMOS2020 catalog also provides estimates of various physical properties derived from spectral energy distribution (SED) fitting, such as photometric redshifts, stellar masses, and star formation rates. Two independent SED-fitting codes, EAZY (Bramme+2008) and LePhare (Ilbert+2006), were employed to estimate photometric redshifts. In this work, we adopt the redshifts derived from LePhare, as they demonstrate better reliability within our magnitude range of interest (see Figure 15 of Weaver+2022). The redshift estimation process utilizes a library of 33 galaxy templates constructed from the stellar population synthesis models of Bruzual+2003 and Ilbert+2009, along with a variety of dust attenuation prescriptions, including the starburst curve from Calzetti+2000, the SMC extinction curve (Prevot+1984), and two modified versions of the Calzetti+2000 law incorporating the 2175 Å bump. The final redshift ($z_{\text{LePh}}$) is defined as the median of the redshift likelihood distribution.
@@
III Analysis for The Unsupervised Method for Morphological Classification
@@
III.2 Convolutional Autoencoder Framework for Noise Reduction
@@
III.3 The APCT Preprocessing Strategies
@@
Figure 5 illustrates the distribution of the mean pixel-value differences before and after a 90 rotation of galaxy images, with and without APCT. The results show that, with APCT, the majority of images exhibit minimal pixel-value differences, whereas without APCT, the differences are significantly larger due to the lack of perfect central symmetry in galaxies. This result demonstrates that the APCT effectively eliminates the effect of rotation on pixel distribution while preserving the core features of the image by converting Cartesian coordinates to polar coordinates. By transforming the problem of rotation invariance into one of translation invariance, the APCT can significantly enhance the rotational classification robustness of CNNs, which will be further validated in  V.1 via more tests.
III.4 Voting strategy and Post-hoc label alignment
@@
IV Analysis for The SML Models and the Experiment Settings
@@ IV.2 Experiment Setting
We further verify that key feature distributions were balanced between the two sets to ensure the model’s generalization capability and the reliability of performance evaluation, as well as to avoid potential biases caused by discrepancies in data distribution. As illustrated in Figure 7, the distributions of galaxy morphological types are consistent between the training and validation sets. Similarly, Figure 8 shows that their distributions in the parameter space of physical properties—such as stellar mass ($M_{\ast}$), redshift ($z$), and other morphological parameters are highly similar (see Section V.3). The Kolmogorov–Smirnov (K–S; Kolmogorov+1933) test results (Table 1) show that all the examined parameters yield $p$-values greater than 0.05, indicating that the training and validation sets follow statistically consistent distributions.
@@
Figure 10: Panels (a) and (b) show the distribution of galaxies in rest-frame NUVrNUV-r vs. rJr-J color space and in the $G-M_{20}$ parameter space, respectively. The contours enclose 70% of the galaxies in each subclass. The scattered points represent 500 galaxies randomly selected from each subclass. Colored diamonds indicate the median values of the distributions, and the associated error bars represent the standard deviations. Panels (c) and (d) display the distributions of Sérsic index ($n$) and effective radius ($r_{\text{e}}$), respectively. In the box plots, the boxes span the interquartile range (IQR) from the first to the third quartile, the central black line marks the median, and the whiskers extend to the minimum and maximum values.
@@ V.1 The effectiveness of CAE and APCT
V.2 Overall morphological classification results and t-SNE Visualization
@@
V.3 Comparisons of Galaxy Properties